OpenStudy (anonymous):

The areas of the rectangles below follow a general pattern. The area of each rectangle is the product of its dimensions, width and length. Identify the pattern among the given areas and then find the area of the rectangle A. Show all steps and justify your observations.

OpenStudy (anonymous):

OK, so there are 3 rectangles && one rectangle is 20, next rec. is 42, next rec. is 72, last rec. is rec. A and i have to find what A is..HELP PLEASE :]

OpenStudy (anonymous):

these numbers are areas of the various rectangles ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, sooo..i divided them by 2 soo i got the L and w's i just Dont know what to d next?

OpenStudy (anonymous):

why did u divide by 2???

OpenStudy (anonymous):

i divided the numbers in the rectangles to get the L and W like 20/2 is 5 * 4 so i assumed that those number's are how i could get the deminsions..

OpenStudy (anonymous):

20/2 is 10

OpenStudy (anonymous):

no it can't be worked tht way... sorry i am also unable to get its solution :(( maybe someone else will help.......

OpenStudy (anonymous):

I'm gonna take a guess that the width by height follows this pattern: 4*5 = 20, 6*7 = 42, 8*9 = 72 and A = 10*11 = 110.

OpenStudy (anonymous):

OR, if A is the first rectangle rather than the last, 2*3 = 6.

OpenStudy (anonymous):

I think estudier is right !!!! good show !!!!

OpenStudy (anonymous):

THANk YOU SOOOOOOOOOOOOOOOOO MUCH estudier !! and thank you harkirat for trying (:

OpenStudy (anonymous):

u r welcome........